(10 ptsFor f(3) = 3+1 a) Find the real part (u) b) Find the imaginary part (v) c) Use u and v to evaluate f(2)atz = 9 - 6
Answer #1

ZA Solution fel zti Z = (x+14) f(z) (utlu) f(2)= ( (atply) f(e) a (xtiy) f2) = ( xtiy) [(x +1) + iy) H2) = (xtiy) [ht) - i4]flie) al er 4= at z = (q-j6) z =(x + 14) Side 9 and y=(-6) x² + y2 of 2) [(992 + (-6)+9) (x+2 +42 (9+992 + (-6)? (81 +36 + 9)